# FM2DS: Few-Shot Multimodal Multihop Data Synthesis with Knowledge Distillation for Question Answering

### v2

@@ Appendix C Investigating FM2DS on Other Models
Table 7: Comparison of model performance across various architectures, sizes, and sample sources (real vs. synthesized by FM2DS). The models were evaluated on 10k samples and the full dataset (23.8k samples for MultiModalQA and 34.2k samples for WebQA). When comparing models tuned on synthesized data with those trained on the full training set, the smallest number of synthetic samples (divisible by 1000) that outperforms models trained on the full datasets is reported. For real sample evaluations, the WebQA training set is used for testing on the WebQA test set, and the same applies to MultiModalQA. Models trained with synthesized samples consistently outperform those trained with equivalent numbers of real samples.
Appendix D Human Evaluation Details
@@
Figure 6: The custom evaluation application used for human evaluation. The application presents each participant with a randomly selected question, relevant Wikipedia pages, and two model-generated answers labeled as Answer A and Answer B. One answer is generated by the pipeline with validation, while the other comes from the pipeline without it. Participants are asked to choose the correct answer and optionally provide feedback on their choice. To minimize bias, the application randomizes the position of each model’s answer.
Upon registration, participants were provided access to a custom evaluation application, as shown in Figure 6. This application was designed to streamline the evaluation process and ensure consistency across participants. For each session, the application randomly selected one of the 100 questions from our human evaluation set of samples. For each question, participants could review the question text, the associated Wikipedia pages, and the generated answers from two methods—one method utilizing the answer validation component and the other without it. To minimize user bias, the application randomly alternated the positioning of the methods’ answers (labeling them as “Answer A” and “Answer B”) so that users could not develop a tendency to select one model over the other based on position alone. After examining the question and relevant Wikipedia content, users were asked to select one of four options to indicate their assessment of answer accuracy: (1) Answer A is correct, (2) Answer B is correct, (3) both answers are correct, or (4) neither answer is correct.
@@
Appendix E M2QA-Bench Samples
Figure 7: Multimodal and multihop reasoning example from M2QA-Bench where the model answers a question about the photograph ”Raising the Flag on Iwo Jima” by synthesizing information from linked documents through a hyperlink, leveraging both visual and tabular data to determine the number of casualties from the USA..
@@ Appendix E M2QA-Bench Samples
Table 9: Performance comparison of different model families fine-tuned on real and synthesized data on M2QA-Bench. The ratios for EM scores and hallucination were calculated from filtered data (EM(Table) refers to the EM score calculated on samples that include the table modality). In the table, $\uparrow$ indicates that larger values are better (all EM values), while $\downarrow$ indicates that smaller values are better (hallucination rate). The scores for EM (Table) and EM (Image) may include samples that also contain other modalities. Larger models and those fine-tuned on synthesized data generally exhibit improved performance with reduced hallucination rates. In MultiModalQA, models demonstrate higher performance on questions involving the image modality compared to those involving tables.
Appendix F Qualitative Analysis
This analysis was conducted for 100 samples from each of the following benchmarks: (1) M2QA-Bench, (2) MultiModalQA, and (3) WebQA. The results are presented in Tables 8, 9, and 10. The responses generated by different models were analyzed across these datasets, focusing on the following metrics:
  1. 1.
    Model accuracy using the exact match (EM) metric.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
@@ Appendix F Qualitative Analysis
Regarding the effect of modalities, results from Tables 8 and 9 suggest that the modalities themselves are not the most critical factor. Instead, the complexity of how the question integrates the modalities plays a more significant role. For M2QA-Bench, models performed better when visual understanding was not required, with tables and text being the primary contributors to the results. In contrast, for MultiModalQA, models tended to perform better on image-based questions, highlighting the importance of how the question leverages the modalities. For questions involving both modalities, smaller models struggled more to produce correct answers, while larger models performed better in terms of EM. It is important to note, however, that due to the substantial difference in the number of samples containing both image and table modalities compared to those with only one modality, the reported results are not directly comparable. Refer to Figures 9 and 10 for the outputs of different model families fine-tuned on either real or synthesized data.
Refer to captionRefer to caption

### v4

@@ Appendix E Investigating FM2DS on Other Models
Table 8: Comparison of model performance across various architectures, sizes, and sample sources (real vs. synthesized by FM2DS). The models were evaluated on 10k samples and the full dataset (23.8k samples for MultiModalQA and 34.2k samples for WebQA). When comparing models tuned on synthesized data with those trained on the full training set, the smallest number of synthetic samples (divisible by 1k) that outperforms models trained on the full datasets is reported. For real sample evaluations, the WebQA training set is used for testing on the WebQA test set, and the same applies to MultiModalQA. Models trained with synthesized samples consistently outperform those trained with equivalent numbers of real samples.
Appendix F Human Evaluation Details
@@
Figure 6: The custom evaluation application was used for human evaluation. The application presents each participant with a randomly selected question, relevant Wikipedia pages, and two model-generated answers labelled as Answer A and Answer B. One answer is generated by the pipeline with validation, while the other comes from the pipeline without it. Participants are asked to choose the correct answer and optionally provide feedback on their choice. To minimize bias, the application randomizes the position of each model’s answer.
After registration, participants were divided into four batches (three participants per batch, each assigned 25 samples, 100 in total) and given access to a custom evaluation app, shown in Figure 6, to review the samples in their assigned batch. This application was designed to streamline the evaluation process and ensure consistency across participants. For each question, participants could review the question text, the associated Wikipedia pages, and the generated answers from two methods—one method utilizing the answer validation component and the other without it. To minimize user bias, the application randomly alternated the positioning of the methods’ answers (labeling them as “Answer A” and “Answer B”) so that users could not develop a tendency to select one model over the other based on position alone. After examining the question and relevant Wikipedia content, users were asked to select one of four options to indicate their assessment of answer accuracy: (1) Answer A is correct, (2) Answer B is correct, (3) both answers are correct, or (4) neither answer is correct.
@@
Figure 7: Distribution of domains in M2QA-Bench.
Figure 8: Distribution of named entities in answers in M2QA-Bench.
Appendix G Additional Statistics & Information on M2QA-Bench
As illustrated in Figure 7, M2QA-Bench encompasses a diverse range of domains. Additionally, the answers span various types of named entities, including people, products, works of art, and more. Figure 8 presents the distribution of named entities found in the answers.
To further examine the diversity of questions in our benchmark—which also reflects the overall characteristics of the data generated by FM2DS—we conducted a 2D t-SNE analysis of question embeddings using ModernBERT (Warner et al., 2024). We sampled 500 questions each from M2QA-Bench, MMQA, WebQA, and ScienceQA (Lu et al., 2022a). ScienceQA serves as a fully human-authored dataset, while MMQA and WebQA primarily use templated questions. As shown in Figure 9, MMQA and WebQA display the least diversity. In contrast, M2QA-Bench, which includes questions generated from FM2DS, demonstrates greater similarity to human-generated data, reflecting a reduced domain gap and improved diversity compared to MMQA and WebQA.
Figure 9: 2D t-sne visualization of ModernBERT embeddings of questions between M2QA-Bench, WebQA, MultimodalQA, and ScienceQA.
Appendix H FM2DS Generalizability to Other Domains
FM2DS can generate domain-specific synthetic data using just a single in-context example and user-provided documents, enabling even small LVLMs to handle domain-specific multimodal multihop QA. As shown in Figure 7, FM2DS’s data (including M2QA-Bench) spans various domains. To further assess its generalizability, we trained InternVL-2-8B on 5k synthesized samples and evaluated it on the health-related dataset PMC-VQA (Zhang et al., 2023) using 50k test samples. We then compared its performance to the pretrained InternVL-2-8B on the same test set. Training improved accuracy from 52.82 to 65.72, demonstrating the benefits of our synthetic data.
Appendix I M2QA-Bench Examples
Figure 10: Multimodal and multihop reasoning example from M2QA-Bench where the model answers a question about the photograph ”Raising the Flag on Iwo Jima” by synthesizing information from linked documents through a hyperlink, leveraging both visual and tabular data to determine the number of casualties from the USA.
@@ Appendix I M2QA-Bench Examples
Appendix J Qualitative Analysis
This analysis was conducted for 100 samples from each of the following benchmarks: (1) M2QA-Bench, (2) MultiModalQA, and (3) WebQA. The results are presented in Tables 9, 10, and 11. The responses generated by different models were analyzed across these datasets, focusing on the following metrics:
  1. 1.
    Model accuracy using the exact match (EM) metric.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
@@
Table 10: Performance comparison of different model families fine-tuned on real and synthesized data on MultimodalQA. The ratios for EM scores and hallucination were calculated from filtered data (EM(Table) refers to the EM score calculated on samples that include the table modality). In the table, $\uparrow$ indicates that larger values are better (all EM values), while $\downarrow$ indicates that smaller values are better (hallucination rate). The scores for EM (Table) and EM (Image) may include samples that also contain other modalities. Larger models and those fine-tuned on synthesized data generally exhibit improved performance with reduced hallucination rates. In MultiModalQA, models demonstrate higher performance on questions involving the image modality compared to those involving tables.
@@
Regarding the effect of modalities, results from Tables 9 and 10 suggest that the modalities themselves are not the most critical factor. Instead, the complexity of how the question integrates the modalities plays a more significant role. For M2QA-Bench, models performed better when visual understanding was not required, with tables and text being the primary contributors to the results. In contrast, for MultiModalQA, models tended to perform better on image-based questions, highlighting the importance of how the question leverages the modalities. For questions involving both modalities, smaller models struggled more to produce correct answers, while larger models performed better in terms of EM. It is important to note, however, that due to the substantial difference in the number of samples containing both image and table modalities compared to those with only one modality, the reported results are not directly comparable. Refer to Figures 13 and 14 for the outputs of different model families fine-tuned on either real or synthesized data. Moreover, Figure 15 shows outputs from different model sizes within the same family, fine-tuned on either real or synthesized data.
Appendix K Synthesizing Data vs. Paraphrasing Exisiting Human Annotated Datasets
Figure 12: hey
Paraphrasing questions from existing datasets introduces surface-level linguistic changes but preserves the original semantic intent and reasoning pathways, offering only marginal improvements in model training. In contrast, the data synthesized by FM2DSis intentionally crafted to introduce diverse question structures, span multiple domains, and require varied types of reasoning, pushing models toward more comprehensive multimodal understanding. To compare these approaches, we trained InternVL-2-8B using 1k samples from three settings: (I) the original MultimodalQA dataset, (II) a paraphrased version of MultimodalQA where questions were reworded using GPT-4o with the prompt ”Please paraphrase the following question: [Question]” and (III) synthesized samples generated by FM2DS. For all conditions, we used full Wikipedia documents as sources. Figure 12 presents the results, showing that while paraphrasing provides a slight improvement, synthesizing new, high-quality samples with FM2DS leads to a substantial performance gain.
Appendix L Statistics on Usages of Each Validation Stage of FM2DS
As described in Section 3 and illustrated in Figure 2, FM2DS incorporates multiple validation stages to enhance data quality. It is essential to analyze how frequently each stage rejects the initially generated outputs. Table 12 presents statistics based on generating 1,000 examples using GPT-4o. Among the stages, question validation has the highest rejection rate, suggesting that this step is the most challenging. This may be because generating a question requires the model to synthesize all relevant knowledge and fully grasp the context. In contrast, answer validation benefits from the guidance provided by the question, making the task relatively easier. Query validation appears to be even more straightforward, as it primarily involves formatting the reasoning steps—something the model has effectively done during answer generation. Additionally, the use of question-specific image captions during answer generation likely contributes to a lower error rate by helping the model locate the correct information only text modality.
Stage Average Rejection Rate Average Rejection Rate
Question Validation 131 11.58%
Answer Validation 76 7.06%
Query Validation 58 5.48%